A loaded penguin sled weighing 80 N rests on a plane inclined at angle θ=20 degress to the horizontal. between the sled and the plane, the coefficent of kinetic of static friction is .25. and the coefficent of kinetic friction is .15 (a) what is the least magnitude of the force F, parallal to the plane what will prevent the sled from slipping down the plane? (b) what is the minimum magnitude F that will start the sled moving up the plane? (c) what value of F is required to move the sled up the plane at constant velocity?
Answer #1
Concepts and reason

The concepts used to solve this question is free body diagram and Newton’s laws of motion.

Initially, draw the free body diagram of the given situation and later apply Newton’s laws of motion to the least force acting on the sled that will prevent it from slipping down the plane. Later determine the minimum force required to move the sled up the plane. At last,calculate the force needed to move the sled up the plane at a constant velocity by using Newton’s laws of motion.

Fundamentals

An object is said to be in motion when it changes its position from one place to another. The object will accelerate if there is a net force acting on the object. The magnitude of the net force is proportional to the mass of the object and acceleration of the object.

The magnitude of the net force acting on the object is,

Fnet=ma{F_{{\rm{net}}}} = ma

Here, mm is the mass of the object and aa is the acceleration of the object.

(a)

The free body diagram for the given situation is drawn below.

Sled
(
(80 N) sin e
(80 N)cos e
80 N

Apply Newton’s laws of motion

The minimum force acting on the sled is,

Fmin=(80N)sinθf{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta - f …… (1)

Here, θ\theta is the angle of inclination and ff is the force of friction.

The force of friction acting on the sled is,

f=μsNf = {\mu _{\rm{s}}}{\rm{N}} …… (2)

Here, μs{\mu _{\rm{s}}} is the coefficient of static friction and NN is the normal force acting on the sled.

The normal force acting on the sled is,

N=(80N)cosθN = \left( {80{\rm{ N}}} \right)\cos \theta …… (3)

Substitute equation (3) in (2).

f=(80N)μscosθf = \left( {80{\rm{ N}}} \right){\mu _{\rm{s}}}\cos \theta …… (4)

Substitute equation (4) in (1).

Fmin=(80N)sinθ(80N)μscosθ{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta - \left( {80{\rm{ N}}} \right){\mu _{\rm{s}}}\cos \theta …… (5)

Substitute 2020^\circ for θ\theta and 0.250.25 for μs{\mu _{\rm{s}}} in the equation (5).

Fmin=(80N)sinθ(80N)μscosθ=(80N)sin20(80N)(0.25)cos20=27.361N18.80N=8.56N\begin{array}{c}\\{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta - \left( {80{\rm{ N}}} \right){\mu _{\rm{s}}}\cos \theta \\\\ = \left( {80{\rm{ N}}} \right)\sin 20^\circ - \left( {80{\rm{ N}}} \right)\left( {0.25} \right)\cos 20^\circ \\\\ = 27.361{\rm{ N}} - 18.80{\rm{ N}}\\\\ = 8.56{\rm{ N}}\\\end{array}

(b)

The diagram for the sled is drawn below.

Motion of sled
Sled
(80 N) sin e
(80 N)cos e
80 N

The minimum force acting on the sled is,

Fmin=(80N)sinθ+f{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta + f …… (6)

Substitute equation (4) in equation (6).

Fmin=(80N)sinθ+(80N)μscosθ{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta + \left( {80{\rm{ N}}} \right){\mu _{\rm{s}}}\cos \theta …… (7)

Substitute 2020^\circ for θ\theta and 0.250.25 for μs{\mu _{\rm{s}}} in the equation (7).

Fmin=(80N)sinθ+(80N)μscosθ=(80N)sin20+(80N)(0.25)cos20=27.361N+18.80N=46.16N\begin{array}{c}\\{F_{\min }} = \left( {80{\rm{ N}}} \right)\sin \theta + \left( {80{\rm{ N}}} \right){\mu _{\rm{s}}}\cos \theta \\\\ = \left( {80{\rm{ N}}} \right)\sin 20^\circ + \left( {80{\rm{ N}}} \right)\left( {0.25} \right)\cos 20^\circ \\\\ = 27.361{\rm{ N}} + 18.80{\rm{ N}}\\\\ = 46.16{\rm{ N}}\\\end{array}

(c)

The net force acting on the sled is,

F=(80N)sinθ+fF = \left( {80{\rm{ N}}} \right)\sin \theta + f …… (8)

The force of friction acting on the sled is,

f=(80N)μkcosθf = \left( {80{\rm{ N}}} \right){\mu _{\rm{k}}}\cos \theta …… (9)

Here, μk{\mu _{\rm{k}}} is the coefficient of kinetic friction.

Substitute equation (9) in equation (8).

F=(80N)sinθ+(80N)μkcosθF = \left( {80{\rm{ N}}} \right)\sin \theta + \left( {80{\rm{ N}}} \right){\mu _{\rm{k}}}\cos \theta …… (10)

Substitute 2020^\circ for θ\theta and 0.150.15 for μk{\mu _{\rm{k}}} in the equation (10).

F=(80N)sinθ+(80N)μkcosθ=(80N)sin20+(80N)(0.15)cos20=27.361N+11.27N=38.63N\begin{array}{c}\\F = \left( {80{\rm{ N}}} \right)\sin \theta + \left( {80{\rm{ N}}} \right){\mu _{\rm{k}}}\cos \theta \\\\ = \left( {80{\rm{ N}}} \right)\sin 20^\circ + \left( {80{\rm{ N}}} \right)\left( {0.15} \right)\cos 20^\circ \\\\ = 27.361{\rm{ N}} + 11.27{\rm{ N}}\\\\ = {\rm{38}}{\rm{.63 N}}\\\end{array}

Ans: Part a

The minimum force needed to prevent the sled from slipping downwards is 8.56N{\bf{8}}{\bf{.56 N}}

Part b

The minimum force needed to prevent the sled from moving up the inclined plane is 46.16N{\bf{46}}{\bf{.16 N}}

Part c

The force required to move the sled up the inclined plane is 38.63N{\bf{38}}{\bf{.63 N}}